Question 3: 1-Suppose that the economy produces two goods tomatoes (T) and cereals (C) What would its PPF look like under conditions of constant opportunity costs? What would it look like with increasing opportunity costs? 2-Using the following data: 20 20 40 60 $1 $2 $8 $8 15 15 12 18 S 5 $10 d. Calculate the countrys nominal GDP levels. For GDP-256S, Pr-4 and Pc-8, Draw the price line.
Answer #1

1. In case of constant opportunity cost the Production Possibility frontier will be straight line. Since, the opportunity cost of producing an additional unit of other good. Hence A straight line.

Good A Good B

While in case of increasing opportunity cost the PPF will be bowed out. Kindly refer the attached picture below

Production Possibility Frontier (PPF)

On one of the axis we have Tomatoes and another Cereals.

2. Nominal GDP = Price of tomato×Quantity of tomato+Price of cereal × Quantity of Cereal

Nominal GDP

Nominal GDP (a) = 5*20 + 1*15 = $ 115

Nominal GDP(b) = 10*20 + 2*15 = $ 230

Nominal GDP(c) = 4*40 + 8*12 = $ 256

Nominal GDP(d) = 4*60 + 8*18 = $ 384

The price line can be determined as follows

First assume 0 unit of tomato is consumed then consumption of cereal = $ 256 /$ 8 = 32 units

When Cereal consumed is zero, then tomato consumed will be equal to $ 256/$4 = 64 units

35 30 r 25 O20 15 O 10 5 0 0 20 40 60 80 Quantity of Tomato

Please contact if having any query thank you.